### v1

@@ 3.1 Interior equilibria
This proposition states that if most consumers reside in region $L$, increasing the freeness of trade will lead to a smooth exodus from region $L$ to region $R$, irrespective of the value of $t(h^{*})$. More integration thus increases the incentives for consumers to distribute more equally among the two regions.
3.2 Logit preferences
For the two region case, and following Tabuchi and Thisse (2002) and Murata (2003), the probability that a consumer will choose to reside in region $L$ is given by the logit model written as:
$P_{L}(h)=\dfrac{e^{u_{L}(h)/\mu}}{e^{u_{L}(h)/\mu}+e^{u_{R}(h)/\mu}},$ (13)
where $\mu\geq 0$ is a scale parameter which measures the dispersion of consumer preferences. If $\mu=0$, consumers do not care about their location preferences but rather solely about relative wages. The number of agents $h$ is the same as the consumer $x$ who is indifferent between living in region $L$ or in region $R$. Therefore, the probability $P_{L}(h)$ is tantamount to the indifferent consumer $x=h$ (Castro et al. 2022). Thus, using (13), we can write:
$h=\dfrac{e^{u_{L}(h)/\mu}}{e^{u_{L}(h)/\mu}+e^{u_{R}(h)/\mu}}.$ (14)
Manipulating (14) yields:
$u_{L}-\mu\ln h=u_{R}-\mu\ln(1-h),$ (15)
which rearranged yields the long-run equilibrium condition, $\Delta u(h)=\Delta t(h),$ where $\Delta t(h)=\mu\left[\ln h-\ln(1-h)\right]$.
Notice from the (15) that $t(h)=\mu\ln h<0$ for $h\in[0,1]$, so that the overall utility here is interpreted as the utility from consumption plus a benefit from living in the most preferred region. The overall utility of a consumer in region $i$ thus lies on the interval $[U_{i},+\infty)$. For a strictly positive $\mu$, the consumer who likes region $R$ the most (at $x=1$) will never want to live in $L$, because, at $h=1$, the overall utility gain from moving from region $L$ to region $R$ is infinite. This means that the Logit model penalizes (benefits) the consumers who are less (more) willing to leave a region very strongly. Therefore, agglomeration in any region is not possible.
This discussion allows us to formalize the following result, which summarizes the possible spatial outcomes under Logit type preferences, depending on the degree of heterogeneity $\mu.$
Proposition 2.
Under Logit type preferences, the spatial distribution depends on the level of consumer heterogeneity as follows:
  • Symmetric dispersion is the only stable equilibrium if:
    $\mu>\mu_{d}\equiv\dfrac{(2\sigma-1)(1-\phi)}{(\sigma-1)(2\sigma+\phi-1)};$ (16)
  • Partial agglomeration $h^{*}\in\left(\tfrac{1}{2},1\right)$ is the only stable equilibrium if $\mu<\mu_{d}$.
Proof.
See Appendix E. ∎
@@ 3.2 Logit preferences
In Figure 3 we show how the spatial distribution of industry changes as regions become more integrated. We set $\sigma=2.5$ and $\mu=0.2$.
[…]
For low levels of the freeness of trade, a unique stable partial agglomeration equilibrium exists where most consumers reside in region $L$. This means that there is just a small fraction of consumers in $R$ that are not willing to forego their preferred region because the gain in consumption goods from doing so is not high enough. As regions become more integrated, the home-market effect becomes weaker and partial agglomeration becomes more symmetric. Finally, for a high level of inter-regional integration, agglomeration forces are so weak that no consumer is willing to leave his most preferred region and only symmetric dispersion is stable.
The preceding analysis is summarized in the two bifurcation diagrams in Figure 4, along a smooth parameter path where $\mu$ and $\phi$ increase. We numerically uncover two supercritical pitchfork bifurcations in $\mu$ and $\phi$. The picture to the right in Figure 4 (increasing $\phi$) differs from other supercritical pitchforks found in the literature (e.g., Pflüger, 2004; Ikeda et al., 2022) crucially in the sense that the direction of change in stability as $\phi$ increases is reversed, i.e., lower trade barriers leads to more symmetric spatial distributions.
Refer to captionRefer to caption
Figure 4: Bifurcation diagrams. To the left, the bifurcation parameter is $\mu\in[0,1]$ with $\phi=0.4$. To the right, we have $\phi\in(0,1)$ as the bifurcation parameter and set $\mu=0.2$. For both scenarios we use $\sigma=2$.
@@ 4 Discussion and concluding remarks
We have seen that heterogeneity in preferences for location alone bears no impact on the relationship between trade integration and spatial inequality, which is a monotonic decreasing one. This contrasts the findings in other works with heterogeneity in consumer preferences, such as Tabuchi and Thisse (2002) and Murata (2003), who show evidence of a bell-shaped relationship between trade integration and spatial inequality. The former’s setting differs from ours because the authors consider an inter-regionally immobile workforce whose role as a dispersive force is enhanced by higher transportation costs.
@@
However, it is particularly worthwhile to discuss the results of Murata (2003). Murata (2003) found that the relationship between trade integration and spatial inequality need not be monotonic and depends on the degree of consumer heterogeneity, which is at odds with our findings. For instance, for an intermediate degree of consumer heterogeneity, Murata finds that increasing trade integration initially fosters agglomeration and later leads to re-dispersion of industry. However, these conclusions can be shown to stem from the author’s particular choice of the value for the elasticity of substitution, $\sigma=1.25$. As we have argued in Section 3, such a low value is empirically implausible. For exceedingly low values $(\sigma<1.71)$, increasing returns at the firm level are too strong. Strong enough that the utility gain at dispersion becomes increasing in $\phi$, instead of decreasing. This would justify an initial concentration of industry as a result of an increase in $\phi.$ However, we have seen that for a plausible range of $\sigma$ the utility gain at dispersion always decreases with $\phi$. Moreover, if $\theta>0$, this holds even for lower values of $\sigma$ [^4] If $\theta\geq 1$, the result holds for $\sigma>1$. Therefore, a higher $\phi$ always promotes more equitable distributions as opposed to asymmetric ones.
@@
Hence, when workers are completely mobile, more trade integration ubiquitously reduces the spatial inequalities between the two regions, irrespective of the degree of heterogeneity in location preferences. Therefore, a de facto lower inter-regional labour mobility induced by consumer heterogeneity alone cannot account for the predictions that a higher inter-regional integration will lead to more unequal spatial development or an otherwise bell-shaped relationship between the two.
@@
By considering that all consumers are allowed to migrate if they so desire, we have shown that a higher inter-regional trade integration always leads to less spatial inequality. This result is independent of the level and impact of consumer heterogeneity. This conclusion may be of potential use for policy makers. Namely, the predictions that globalization is likely to lead to more spatial inequality (World Bank, 2009) in the future may be reversed if policies are undertaken to promote inter-regional mobility.
@@
This derivative is positive for $w\in[\phi^{1/\sigma},\phi^{-1/\sigma}]$, which implies that $w(h)$ is increasing in $[0,1]$.
@@
which is positive if $h>1/2$, because the latter implies $w>1$ (because $w$ is increasing in $h$). As a result, we have $dw/d\phi<0$. Therefore, the nominal wage $w$ is decreasing in the freeness of trade when $L$ is the largest region ($h>1/2$).
@@
Proof of Lemma 1
Proof.
Taking the derivative of (10) with respect to $h$ yields:
$\dfrac{d\Delta u}{dh}=-\frac{w^{-(\theta+\sigma)}\left\{w^{\theta}\left(h\phi w^{1-\sigma}-h+1\right)^{-\frac{\theta+\sigma-2}{\sigma-1}}\left[-h(\sigma-1)\phi w^{\prime}-w^{\sigma}+\phi w\right]+\left(hw^{1-\sigma}-h\phi+\phi\right)^{-\left(1+\frac{\theta}{\sigma-1}\right)}\varOmega\right\}}{(\sigma-1)},$ (19)
where:
$\Omega=-w\left(hw^{1-\sigma}-h\phi+\phi\right)^{\frac{1}{\sigma-1}}\left[-h(\sigma-1)w^{\prime}-\phi w^{\sigma}+w\right]-(\sigma-1)w^{\sigma}w^{\prime}\left(hw^{1-\sigma}-h\phi+\phi\right)^{\frac{\sigma}{\sigma-1}},$
Substituting for $w^{\prime}$ using (18) from Appendix A, and given that any partial equilibrium $h^{*}$ must satisfy the short-run equilibrium condition given by (8), we reach:
$\left.\dfrac{d\Delta u}{dh}\right|_{h=h^{*}}=\zeta\left\{\varphi\left[\dfrac{w^{\sigma-1}+w^{-\sigma}-(1+w^{-1})\phi}{1-\phi^{2}}\right]^{\frac{\theta-1}{\sigma-1}}-\psi\left[\dfrac{w^{2\sigma}-(w+1)\phi w^{\sigma}+w}{w\left(1-\phi^{2}\right)}\right]^{\frac{\theta+\sigma-2}{\sigma-1}}\right\},$ (20)
@@
$\zeta$ $=\dfrac{w^{-(\theta+\sigma+1)}}{(\sigma-1)\left[(\sigma-1)\phi+(\sigma-1)\phi w^{2\sigma}+\left(-2\sigma+\phi^{2}+1\right)w^{\sigma}\right]};$
$\varphi$ $=w\left\{\phi\left[\sigma+(\sigma-1)w\right]w^{\sigma}-2\sigma w+w\right\}\left[w^{2\sigma}-(w+1)\phi w^{\sigma}+w\right];$
$\psi$ $=\left(1-\phi^{2}\right)w^{\theta+\sigma+1}\left[(2\sigma-1)w^{\sigma}-\sigma(w+1)\phi+\phi\right].$
Since an interior equilibrium is stable if $d\Delta u/dh<d\left[t(h)-t(1-h)\right]/dh$ at $h=h^{*}$, it is stable if (20) is negative, which concludes the proof. ∎
@@
Assume that partial agglomeration $h^{*}=h\in\left(\frac{1}{2},1\right)$ exists and is stable. Then, according to Castro et al. (2020, Prop. 9, pp. 197), it becomes more symmetric if $\frac{d\Delta u}{d\phi}<0$. We have:
@@
$-\frac{w}{a_{3}}\left\{a_{1}\left[-\frac{\left(\phi^{2}-1\right)w^{\sigma+1}}{w^{2\sigma}-(w+1)\phi w^{\sigma}+w}\right]^{-\frac{\theta+\sigma-2}{\sigma-1}}+a_{2}\left[\frac{w-w\phi^{2}}{w^{2\sigma}-(w+1)\phi w^{\sigma}+w}\right]^{-\frac{\theta+\sigma-2}{\sigma-1}}\right\}$ (21)
@@
$a_{3}$ $=(\sigma-1)\left[w^{1-\sigma}+w^{\sigma}-(w+1)\phi\right]\left[(\sigma-1)\phi+(\sigma-1)\phi w^{2\sigma}+\left(-2\sigma+\phi^{2}+1\right)w^{\sigma}\right]$
@@
$-\frac{w\left[\frac{w-w\phi^{2}}{w^{2\sigma}-(w+1)\phi w^{\sigma}+w}\right]^{-\frac{\theta+\sigma-2}{\sigma-1}}\left(a_{1}w^{-\sigma\frac{\theta+\sigma-2}{\sigma-1}}+a_{2}\right)}{a_{3}}.$
@@
$-\frac{a_{1}w^{-\sigma\frac{\theta+\sigma-2}{\sigma-1}}+a_{2}}{a_{3}}.$
@@
Next, evaluating the numerator at $\theta=1$, we get:
$w^{-\sigma}\left(\phi w^{\sigma}-1\right)\left[-2\sigma+2(\sigma-1)\phi w^{\sigma}+\phi^{2}+1\right]+w^{-\sigma}\left(w^{\sigma}-\phi\right)\left[2(\sigma-1)\phi+\left(-2\sigma+\phi^{2}+1\right)w^{\sigma}\right]<0.$
Differentiating the numerator with respect to $\theta$ yields:
$\frac{\ln(w)w^{\frac{\theta-1}{\sigma-1}+\sigma}\left(\phi w^{\sigma}-1\right)\left[-2\sigma+2(\sigma-1)\phi w^{\sigma}+\phi^{2}+1\right]}{\sigma-1},$
which is positive because $1<w^{\sigma}<1/\phi$ and $\left[-2\sigma+2(\sigma-1)\phi w^{\sigma}+\phi^{2}+1\right]<0$. Thus, the numerator of (21) remains negative for all $\theta<1$. We conclude that (21) is negative and thus $\frac{d\Delta u}{d\phi}<0$ for any $\theta\in[0,1]$. This concludes the proof. $\square$
@@
Replacing in (20) and rearranging, partial agglomeration is stable if:
@@
$\mu>\mu_{d}\equiv\dfrac{(2\sigma-1)(1-\phi)}{(\sigma-1)(2\sigma+\phi-1)}.$
@@
We can also see that $\mu_{p}$ approaches zero as $w^{\sigma}$ approaches $\phi^{-1}$. Therefore, we conclude that $0<\mu_{p}<\mu_{d}$. Under the assumption that only one partial agglomeration equilibrium $h^{*}\in(1/2,1)$ exists, we have two possibilities: (i) if $\mu\in(\mu_{p},u_{d})$, partial agglomeration is the only stable equilibrium; and (ii) if $\mu>\mu_{d},$ symmetric dispersion is the only stable equilibrium. Assume now, by way of contradiction, that $\mu\in(0,\mu_{p})$. Then both dispersion and partial agglomeration are unstable. Since the state space is one dimensional, this would require agglomeration to be stable. However, we know that agglomeration is always unstable, which implies that either partial agglomeration or dispersion are stable. Hence, $\mu\notin(0,\mu_{p})$. Therefore, symmetric dispersion is the only stable equilibrium if $\mu>\mu_{d}$; otherwise, partial agglomeration is the only stable equilibrium. $\square$

### v2

@@ 3.2 Logit preferences for location
Figure 3: Utility differential $\Delta u$ (dashed lines) and penalty differential $t(h)-t(1-h)$ (thick line) for increasing levels of $\phi$. For $\phi=0.3$ (upper dashed line), dispersion is unstable and partial agglomeration is stable. For $\phi=0.5$ (medium dashed line) dispersion is unstable, partial agglomeration is stable and is less asymmetric. For $\phi=0.9$ (lower dashed line) dispersion is stable and is the only equilibrium. The values for the parameters are $\sigma=2.5$, $\theta=0$ and $\mu=0.2$.
In Figure 3 we show how the spatial distribution of industry changes as regions become more integrated. We set $\sigma=2.5$ and $\mu=0.2$. For low levels of the freeness of trade, a unique stable partial agglomeration equilibrium exists where most consumers reside in region $L$. This means that there is just a small fraction of consumers in $R$ that are not willing to forego their preferred region because the gain in consumption goods from doing so is not high enough. As regions become more integrated, the home-market effect becomes weaker and partial agglomeration becomes more symmetric. Finally, for a high level of inter-regional integration, agglomeration forces are so weak that no consumer is willing to leave his most preferred region and only symmetric dispersion is stable.
Figure 4: Bifurcation diagrams. To the left, the bifurcation parameter is $\mu\in[0,1]$ with $\phi=0.4$. To the right, we have $\phi\in(0,1)$ as the bifurcation parameter and set $\mu=0.2$. For both scenarios we use $\sigma=2$ and $\theta=0$.
The preceding analysis is summarized by the two bifurcation diagrams in Figure 4, along a smooth parameter path where $\mu$ (left picture) and $\phi$ (right picture) increase. We uncover two pitchfork bifurcations in $\mu$ and $\phi$. We confirm analytically that they are supercritical for the case of logit preferences and under Assumption 1.777Employing $\phi$ as the bifurcation parameter, we have $\tfrac{\partial^{3}\Delta V}{\partial h^{3}}(\tfrac{1}{2};\phi_{b})<0$, with $\phi_{b}=\frac{(2\sigma-1)\left[\mu(\sigma-1)-1\right]}{-(\mu+2)\sigma+\mu+1}$ and $\sigma\in\left(2,\frac{\mu+1}{\mu}\right)$. Then by the proof of Lemma 3 (Appendix B) the pitchfork is supercritical. For the case of $\mu$ as the bifurcation parameter, see the proof of Proposition 3 (Appendix C). This implies that a curve of stable partial agglomeration equilibria branches in the direction of $\mu<\mu_{d}$ (left picture) and $\phi<\phi_{b}$ (right picture), showing that lower trade barriers or a higher heterogeneity promote more even spatial distributions. The picture to the right differs from other supercritical pitchforks found in the literature (e.g., Pflüger, 2004; Ikeda et al., 2022) crucially in the sense that the direction of change in stability as $\phi$ increases is reversed, i.e., lower trade barriers leads to more symmetric spatial distributions, as expected from our analysis. Regarding the effect of $\mu$, it is quite intuitive that a higher scale of heterogeneity increases local dispersion forces (i.e., dispersion forces that do not depend on $\phi$) and thus also promotes a more even distribution of economic activities.
@@ 4 Discussion and concluding remarks
We have seen that heterogeneity in preferences for location alone bears no impact on the relationship between trade integration and agglomeration, which is a monotonic decreasing one. This contrasts the findings in other works with heterogeneity in consumer preferences, such as Tabuchi and Thisse (2002) and Murata (2003), who show evidence of a bell-shaped relationship between trade integration and the spatial distribution of industry, whereby agglomeration but is followed by a re-dispersion phase as trade integration increases. The former’s setting differs from ours because the authors consider an inter-regionally immobile workforce whose role as a dispersive force is enhanced by higher transportation costs.
@@
However, it is particularly worthwhile to discuss the results of Murata (2003), who uses a model that is a particular case of ours if we set $\theta=0$ and consider logit type heterogeneity. Murata (2003) found that the relationship between trade integration and agglomeration need not be monotonic and depends on the degree of consumer heterogeneity, which is at odds with our findings. For instance, for an intermediate degree of consumer heterogeneity $(\mu)$, Murata finds that increasing trade integration initially fosters agglomeration and later leads to re-dispersion of industry. However, these conclusions can be shown to stem from the author’s particular choice of the value for the elasticity of substitution, $\sigma=1.25$. As we have argued in Section 3, such a low value is empirically implausible. For exceedingly low values $(\sigma<1.71)$, increasing returns at the firm level are too strong. Strong enough that the utility gain at dispersion becomes increasing in $\phi$, instead of decreasing. This would justify an initial concentration of industry as a result of an increase in $\phi.$ However, we have seen that for a plausible range of $\sigma$ the utility gain from consumption at symmetric dispersion always decreases with $\phi$. Moreover, if $\theta>0$, this holds even for lower values of $\sigma$ [^8] If $\theta\geq 1$, the result holds for $\sigma>1$. Therefore, a higher $\phi$ always promotes more equitable distributions as opposed to asymmetric ones.
@@
Hence, when workers are completely mobile, more trade integration ubiquitously leads to a more even dispersion of spatial distributions among the two regions, irrespective of the degree of heterogeneity in location preferences. Therefore, a de facto lower inter-regional labour mobility induced by consumer heterogeneity alone cannot account for the predictions that a higher inter-regional integration will lead to more spatial inequality or an otherwise bell-shaped relationship between the two.
@@
By considering that all consumers are allowed to migrate if they so desire, we have shown that a higher inter-regional trade integration always leads to more dispersed spatial distributions. This result is independent of the level and impact of consumer heterogeneity. This conclusion may be of potential use for policy makers. Namely, the predictions that globalization is likely to lead to higher spatial inequality in the future (World Bank, 2009) may be reversed if policies are undertaken to promote inter-regional mobility.
@@
This derivative is positive for $w\in[\phi^{1/\sigma},\phi^{-1/\sigma}]$, which implies that $w(h)$ is increasing in $[0,1]$. This concludes the proof for the first statement.
@@
which is positive if $h>1/2$, because the latter implies $w>1$ (because $w$ is increasing in $h$). As a result, we have $dw/d\phi<0$. Therefore, the nominal wage $w$ is decreasing in the freeness of trade when $L$ is the largest region ($h>1/2$). Conversely, we have $dw/d\phi>0$ if $h>1/2$. This concludes the proof of the second statement. $\square$
@@
Proof of Lemma 2:
Taking the derivative of (10) with respect to $h$, substituting for $w^{\prime}$ using (18) from Appendix A, and evaluating at $h^{*}$ using the short-run equilibrium condition given by (8), we reach:
$\left.\dfrac{d\Delta u}{dh}\right|_{h=h^{*}}=\zeta\left\{\left(\varphi w^{\sigma\frac{1-\theta}{\sigma-1}}+\frac{\psi}{w}\right)\left\{\frac{\left(1-\phi^{2}\right)w}{w^{2\sigma}-\left[\varphi(w+1)w^{\sigma}\right]+w}\right\}^{\frac{1-\theta}{\sigma-1}}\right\},$ (19)
@@
$\zeta$ $=\dfrac{w^{2\sigma}-\left[\phi(w+1)w^{\sigma}\right]+w}{(\sigma-1)\left[(\sigma-1)\phi+(\sigma-1)\phi w^{2\sigma}+\left(-2\sigma+\phi^{2}+1\right)w^{\sigma}\right]};$
$\varphi$ $=w^{-\theta-\sigma}\left\{\phi\left[\sigma+(\sigma-1)w\right]w^{\sigma}-2\sigma w+w\right\};$
$\psi$ $=(\sigma-1)\phi+(1-2\sigma)w^{\sigma}+\sigma\phi w.$
The fact that an interior equilibrium is stable if $d\Delta u/dh<d\left[t(h)-t(1-h)\right]/dh$ at $h=h^{*}$ concludes the proof. $\square$
Proof of Lemma 3:
Suppose there exists a value of $\phi=\phi_{b}\in(0,1)$ such that $\dfrac{d\Delta u}{dh}\left(\tfrac{1}{2};\phi_{b}\right)=t^{\prime}\left(\tfrac{1}{2};\phi_{b}\right)$. Let $\Delta V=V_{L}-V_{R}$. Then it is possible to show the following:
$\frac{\partial\Delta V}{\partial h}\left(\tfrac{1}{2};\phi_{b}\right)$ $=0;\ \ \frac{\partial^{2}\Delta V}{\partial h^{2}}\left(\tfrac{1}{2};\phi_{b}\right)=0;\ \ \frac{\partial\Delta V}{\partial\phi}\left(\tfrac{1}{2};\phi_{b}\right)=0;$
$\frac{\partial^{2}\Delta V}{\partial h\partial\phi}\left(\tfrac{1}{2};\phi_{b}\right)$ $=-\frac{4\eta\left(\frac{1+\phi_{b}}{2}\right)^{\frac{1-\theta}{\sigma-1}}\left[\theta(1-\phi_{b})+2\sigma+\phi_{b}-3\right]}{(\sigma-1)^{2}(\phi_{b}+1)^{2}};$
The last expression is negative if we assume that $\sigma>1+\sqrt{2}/2$. Next, we have:
$\frac{\partial^{3}\Delta V}{\partial h^{3}}\left(\tfrac{1}{2};\phi_{b}\right)=\frac{16(\sigma-2)(2\sigma-3)(1-\phi_{b})^{3}\left(\frac{1+\phi_{b}}{2}\right)^{\frac{1}{\sigma-1}}}{(\sigma-1)^{3}(\phi_{b}+1)^{3}}$ $-2t^{\prime\prime\prime}\left(\frac{1}{2}\right).$
If $\tfrac{\partial^{3}\Delta V}{\partial h^{3}}\neq 0$, then, according to Guckenheimer and Holmes (2002), the model undergoes a pitchfork bifurcation at symmetric dispersion and a curve of partial agglomeration equilibria branches at $\phi=\phi_{b}.$ If the pitchfork is supercritical ($\tfrac{\partial^{3}\Delta V}{\partial h^{3}}<0$), the curve branches in the direction of $\phi<\phi_{b}$ and partial agglomeration is stable. If it is subcritical ($\tfrac{\partial^{3}\Delta V}{\partial h^{3}}>0)$, it branches in the direction of $\phi>\phi_{b}$ and partial agglomeration is unstable. The criticality of the bifurcation depends on the specific functional form of $t(x)$. This concludes the proof. $\square$
@@
Assume that partial agglomeration $h^{*}=h\in\left(\frac{1}{2},1\right)$ exists and is stable. Then, according to Castro et al. (2022, Prop. 9, pp. 197), it becomes more symmetric if $\frac{d\Delta u}{d\phi}<0$. We have:
@@
$-\frac{w}{a_{3}}\left\{a_{1}\left[\frac{\left(1-\phi^{2}\right)w^{\sigma+1}}{w^{2\sigma}-(w+1)\phi w^{\sigma}+w}\right]^{-\frac{\theta+\sigma-2}{\sigma-1}}+a_{2}\left[\frac{w(1-\phi^{2})}{w^{2\sigma}-(w+1)\phi w^{\sigma}+w}\right]^{-\frac{\theta+\sigma-2}{\sigma-1}}\right\}$ (20)
@@
$a_{3}$ $=(\sigma-1)\left[w^{1-\sigma}+w^{\sigma}-(w+1)\phi\right]\left[(\sigma-1)\phi+(\sigma-1)\phi w^{2\sigma}+\left(-2\sigma+\phi^{2}+1\right)w^{\sigma}\right].$
@@
$-\frac{w\left[\frac{w(1-\phi^{2})}{w^{2\sigma}-(w+1)\phi w^{\sigma}+w}\right]^{-\frac{\theta+\sigma-2}{\sigma-1}}\left(a_{1}w^{-\sigma\frac{\theta+\sigma-2}{\sigma-1}}+a_{2}\right)}{a_{3}}.$
@@
$\Psi\equiv-\frac{a_{1}w^{-\sigma\frac{\theta+\sigma-2}{\sigma-1}}+a_{2}}{a_{3}}.$
@@
Next, taking the limit of $\Psi$ as $\theta\rightarrow+\infty$, we get:
$\lim_{\theta\rightarrow+\infty}\Psi=w^{-\sigma}\left(w^{\sigma}-\phi\right)\left[2(\sigma-1)\phi+\left(-2\sigma+\phi^{2}+1\right)w^{\sigma}\right]<0.$
Differentiating $\Psi$ with respect to $\theta$ yields:
$\frac{\ln(w)w^{-\frac{\sigma(\theta+\sigma-2)}{\sigma-1}-\theta+1}\left(\phi w^{\sigma}-1\right)\left[-2\sigma+2(\sigma-1)\phi w^{\sigma}+\phi^{2}+1\right]}{\sigma-1},$
which is positive because $1<w^{\sigma}<1/\phi$ and $\left[-2\sigma+2(\sigma-1)\phi w^{\sigma}+\phi^{2}+1\right]<0$. Thus, $\Psi$ remains negative for all $\theta\geq 0$. We conclude that (20) is negative and thus $\frac{d\Delta u}{d\phi}<0$ for any $\theta\in\left[0,1\right)\cup\left(1,+\infty\right)$. For $\theta=1$, take $u_{i}=\ln(C_{i})$ to verify that the derivative is negative. We conclude that $\frac{d\Delta u}{d\phi}<0,\forall\theta\geq 0$. This finishes the proof. $\square$
@@
Using $\theta=1$ in $\tfrac{d\Delta u}{dh}$ given by (19), and rearranging, partial agglomeration is stable if:
@@
$\mu>\mu_{d}\equiv\frac{(2\sigma-1)(1-\phi)}{(\sigma-1)(2\sigma+\phi-1)}.$
@@
We can also see that $\mu_{p}$ approaches zero as $w^{\sigma}$ approaches $\phi^{-1}$. Therefore, we conclude that $0<\mu_{p}<\mu_{d}$. We can show that symmetric dispersion undergoes a pitchfork bifurcation at $\mu=\mu_{d}$, following a procedure similar to that of the proof of Lemma 3 (see Appendix B) by replacing $\phi$ with $\mu$ as the bifurcation parameter. The corresponding derivatives have the same signs when evaluated at $\mu=\mu_{d}$. Furthermore, we have:
$\dfrac{\partial^{3}\Delta V}{\partial h^{3}}(\tfrac{1}{2};\mu_{d})=-\frac{64(1-\phi)\phi\left\{\sigma\left[\phi(\phi+2)+5\right]-\phi^{2}-3\right\}}{(\sigma-1)(\phi+1)^{3}(2\sigma+\phi-1)}<0,$
which ensures that the pitchfork is supercritical and proves that a branch of stable partial equilibria emerges in the direction of $\mu<\mu_{d}$. Since the state space is one-dimensional, multiple partial agglomeration equilibria require the interchange of stability between consecutive equilibria, which cannot happen when $\mu\in(\mu_{p},\mu_{d})$. [^9] This holds even if equilibria are irregular. We thus have two possibilities: (i) if $\mu\in(\mu_{p},u_{d})$, partial agglomeration is the only stable equilibrium and is unique; and (ii) if $\mu>\mu_{d},$ symmetric dispersion is the only stable equilibrium.
Next, we know from the text that a partial agglomeration equilibrium always exists if $\mu<\mu_{d}$. Assume now, by way of contradiction, that $\mu\in(0,\mu_{p})$. Then both dispersion and partial agglomeration are unstable. Since the state space is one-dimensional, there can be only one unstable partial agglomeration, which would require agglomeration to be stable. However, we know that agglomeration is always unstable, which implies that partial agglomeration is always stable. Hence, $\mu\notin(0,\mu_{p})$. Therefore, symmetric dispersion is the only stable equilibrium if $\mu>\mu_{d}$; otherwise, partial agglomeration is the only stable equilibrium. $\square$